OpenStudy (anonymous):

limit x->1 x^3 -3x +2 / x-1

OpenStudy (anonymous):

I get 0/0 and o supose its wrong :(

OpenStudy (anonymous):

do you know lhospital's law?

OpenStudy (anonymous):

I have read about it I dont know how to use it

OpenStudy (anonymous):

you need to build the derivats of the two functions

OpenStudy (anonymous):

you take the derivative of the nominator and the denominator (x^3-3x+2)' = ? (x-1)' = ?

OpenStudy (anonymous):

If i set 1 in for x I get 1^3 - 3 (1) +2 / 1-1 and get 0/0 = 0

OpenStudy (anonymous):

limit is x to 1

OpenStudy (anonymous):

(x^3-3x+2)' = 3x^2 -3 (x-1)' = 1 now : lim x->1 of (3x^2 - 3) / 1 = 0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\] x^3 - 3x +2 / x -1

OpenStudy (anonymous):

okay so It does not have to be 1? since the limit goes to 1

OpenStudy (anonymous):

it goes to 0 as coolsector wrote

OpenStudy (anonymous):

if you dont know lhospitals we can do as well : nominator : x^3 -x -2x -2 x(x^2 - 1) -2(x-1) x(x-1)(x+1) -2(x-1) so the limit : (x-1)(x(x+1)-2) / (x-1) -> lim of x(x+1)-2 as x goes to 1 = 0.

OpenStudy (anonymous):

this is another way if you dont know lhospital's..

OpenStudy (anonymous):

Okay Thanks :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

by l'Hospital's rule differentiating the numerator and denominator... we get Lim x->1 3x^2-3/1 applying limit we can get 3(1)^2 -3 = 3-3 =0